OpenStudy (anonymous):

x - 3y + z = -15 2z + y - z =-2 x + y + 2z = 1 Solve using substitution to find the ordered triple. Help me work through this. I don't know what I'm doing.

OpenStudy (anonymous):

@zepdrix @SolomonZelman

OpenStudy (wolf1728):

are you sue the second equation is typed correctly? 2z + y - z =-2

OpenStudy (anonymous):

* 2x + y - z = -2

OpenStudy (help!!!!):

use eliminationn

OpenStudy (anonymous):

I know that

OpenStudy (anonymous):

I've worked out most of it but I messed up along the way and can't figure out where I went wrong

OpenStudy (help!!!!):

so you elimination for the first and second and the second to the third one

OpenStudy (anonymous):

That's not how my teacher worked it

OpenStudy (help!!!!):

well here the real way lol

OpenStudy (anonymous):

That doesn't help. I need to actually see the equations being worked

OpenStudy (help!!!!):

how much do you know?

OpenStudy (anonymous):

I don't know what I know? I know I use elimination on the first and second, then first and third, then put those equations together and find the variables from there

OpenStudy (anonymous):

Thank you for the link, I'll check it out. But I need someone to work it out with me so I can understand what I'm doing on the test. @wolf1728

OpenStudy (wolf1728):

Multiply the second equation by -1 x - 3y + z = -15 2x + y - z =-2 x + y + 2z = 1 giving us x - 3y + z = -15 -2x - y + z =2 x + y + 2z = 1 Now we can add all 3 equations -3y +4z = -12

OpenStudy (anonymous):

My teacher doesn't do it that way. I don't know if he's teaching a different method or what

OpenStudy (wolf1728):

Okay, how does the teacher do it?

OpenStudy (anonymous):

He uses elim. on the first and second equations. Then repeats that on the first and third. Then I add the two new equations to find my variable. Then I plug that variable back into an original equation to find the other two variables

OpenStudy (anonymous):

Then I make an ordered triple like (2, 3, 4)

OpenStudy (shoba):

do u want to to solve in matrix?

OpenStudy (anonymous):

What does that mean?

OpenStudy (wolf1728):

shoba - you mean determinants?

OpenStudy (help!!!!):

pssst seriously google it or watch a youtube video. It's more reliable than us

OpenStudy (help!!!!):

lol

OpenStudy (anonymous):

I don't learn that way. I need. somebody. working. with. meeeeeee...

OpenStudy (help!!!!):

yeah youtube video work with you step by step

OpenStudy (help!!!!):

use khan academy

OpenStudy (wolf1728):

How about solving for x in equation 1 then substituting that value of x in equations 2 and 3?

OpenStudy (anonymous):

Because I'm not allowed to do that on the test. I'm in public school and I have 40 something minutes on my test and I have to work my problems out with the method my teacher teaches me with. I can't just decide to change his choice of method

OpenStudy (shoba):

so how u actually want to solve this question? elimination or substitution

OpenStudy (wolf1728):

Using elimination on the first and second equations then using that on the first and third? Anyone know how to do this?

OpenStudy (wolf1728):

(That's the method babysmith's teacher uses).

OpenStudy (shoba):

yeah i know.... x - 3y + z = -15----1 2x + y -z = 2------2 x + y + 2z = 1------3 1 + 2 x + 2x -3y +y +z - z = -15 +2 3x +4y = -13 -----4 from equation 4----- x = \[\frac{ -4y -13 }{ 3 }\] the subtitude into equation 1 and 3 \[\frac{ -4y -13 }{ 3 }\] - 3y + z = =-15 \[\frac{ -4y -13 }{ 3 }\] + y +2z = 1

OpenStudy (shoba):

after the subtituition we will get multiply both with 3 (-4y +13/3) + y -z = 2 -4y -13 + 3y -3z = 6 -y -3z = 19--------5 (-4y -13/3) + y + 2z = 1 -4y - 13 + 3y + 6z = 3 -y + 6z =16--------6 equation 5 - 6 -3z -6z = 19 - 16 -3z = 3 z = -1 from equation 5 ----- y =-3z -19 = -3(-1) -16 = -13 from equation 1----- x = -15 + 3y + z = 5 + 3(-13) +(-1) = -55

OpenStudy (anonymous):

You don't have to solve it anymore. I can't use that method

OpenStudy (wolf1728):

CORRECTION TO CALCULATIONS A) x - 3y + z = -15 B) 2x + y - z =-2 C) x + y + 2z = 1 in equation A solve for X A) x = 3y -z -15 substitute this into equation B B) 2 * (3y -z -15) + y -z = -2 B) 6y -2z -30 + y -z = -2 B) 7y -3z =28 substitute this into equation C C) (3y -z -15) +y + 2z =1 C) 4y +z = 14

OpenStudy (shoba):

my answer is wrong the answer is x = -3,y=4 and z = 0

OpenStudy (wolf1728):

Yes, I have a calculator here: http://www.1728.org/unknwn3.htm the answers are: X=-3 Y=4 Z=0